OpenStudy (anonymous):

please answer my question If θ is in quadrant 1 and tan θ = 5/12. The value of tan 4θ/5 = ...

OpenStudy (anonymous):

does this ask for \(\tan(\frac{4\theta}{5})\)

OpenStudy (anonymous):

yeah.. hows the answer? help pls

OpenStudy (anonymous):

i have to think, because i don't know of a formula for \(\tan(\frac{ax}{b})\) give me a minute

OpenStudy (anonymous):

oh please :)

OpenStudy (anonymous):

are you sure this is the exact question: if \(\tan(\theta)=\frac{5}{12}\) then find \(\tan(\frac{4\theta}{5})\)

OpenStudy (anonymous):

this looks like a set up for a \(5-12-13\) right triangle, but i don't see how to use it here

OpenStudy (anonymous):

tan(4θ/5) and tan 4θ/5. are there same question? my teacher wrote tan 4θ/5 not tan (4θ/5)

OpenStudy (anonymous):

yeah it is the same

OpenStudy (anonymous):

the best i can do is whip out a calculator and compute \[\tan\left(\frac{4\tan^{-1}(\frac{5}{12})}{5}\right)\]

OpenStudy (anonymous):

if i answer: tan [4x(5/12)]/5 thats false?

OpenStudy (anonymous):

yes, that is false

OpenStudy (anonymous):

how can the ancwer like that? im confuse :(